OpenStudy (anonymous):

Whats the distance from the hiker to the tree when the tree is 83.3 m tall and the angle of elevation from where the hiker is standing is 26 degrees?

OpenStudy (rane):

isn't it physics?

OpenStudy (anonymous):

Nope, its triangles D:

OpenStudy (anonymous):

 So, tan 26 = 83.3/distance.

OpenStudy (anonymous):

I multiply Distance to both side then divide by Tan26 right?

OpenStudy (rane):

The tangent of the angle is the opposite side over the near side. opposite is the height, near side is the distance. so tan 26 = 83.3/x x= 83.3/ (tan 26) just solve now

OpenStudy (anonymous):

^ That's precisely what I said LOL